OpenStudy (anonymous):

a student takes out two loans totaling $16,000 to help pay for college expenses. One loan is at 5% simple interest, and the other is at 6% simple interest. the first-year interest is $890. Find the amount of the loan at 6%.

OpenStudy (anonymous):

x+y=16000 0.05x+0.06x=890

OpenStudy (anonymous):

you can find answer from these.

OpenStudy (slaaibak):

Suppose the loan at 6% is x Therefore, the other loan at 5% is 16 000 - x Now, the interest of both loans would equal 890. Therefore: x* .06 + (16000 - x) * .05 = 890 .01x = 890 - 800 x = 9000 Therefore the first loan = 9000 @ 6%

OpenStudy (anonymous):

Thanks so much guys!